# DeepScholar-Bench: A Live Benchmark and Automated Evaluation for Generative Research Synthesis

### v1

@@ 5.4. Human Agreement Study
Finally, we study how well our LLM-based evaluation aligns with human judgments, a critical question to validate effectiveness of our fully automated evaluation approach. Overall, we find that each of the metrics we introduce for the DeepScholar-bench task exhibit high agreement between LLM-based judgments and human annotations. We collect over 200 expert annotations, and Table 4 shows the agreement score between human and LLM labelers for organization pairwise comparisons, nugget importance labels, graded relevance scores and reference importance labels. Overall, the results demonstrate above agreement scores across each, reflecting the effectiveness of the DeepScholar-bench evaluation approach.
6. Related Work
Long-form Synthesis Benchmarks. Several recent benchmarks offer datasets designed to evaluate long-form research synthesis tasks, however, their design often involves manually-curated queries, which are prone to data staleness, contamination and limited scalability. DeepScholar-bench addresses these challenges by providing an automated data pipeline and live benchmark with realistic, challenging and recent research synthesis tasks, in contrast to prior works. Specifically, ScholarQABench (Asai et al., 2024) creates realistic literature review questions with detailed answer rubrics, written by expert PhD annotators from the computer science, biomedicine, physics and neuroscience domains. Similarly, OpenResearcher (Zheng et al., 2024) constructs a dataset of around 38 scientific text summarization queries, among other research-style questions, by recruiting experts to write and annotate queries. Likewise, DeepConsult (noa, [n.d.]k) provides a suite of expert-curated, open-ended research queries related to business and consulting. Unfortunately, these expert-curated benchmarks, are expensive to construct, and difficult to update, causing them to quickly become outdated, as new information becomes available, These prior benchmark datasets also risk data contamination, as new models are trained on publically available data. Our work instead proposes a scalable, automated pipeline to provide a live, evergreen dataset that reflects diverse and recent research queries.
Similarly to our approach, the FreshWiki dataset (Shao et al., 2024) is constructed using an automated data pipeline; however, the dataset task focuses on the generation of Wikipedia-like articles, whereas this work focuses on a significantly different and difficult synthesis task based on cutting-edge research. The FreshWiki framework focuses on evaluating the article pre-writing stage as well as the generated full-length article based on a ground truth Wikipedia article and established criteria of a good Wikipedia article (noa, 2025d). This task reflects an interesting area of study; however, in this work, we instead focus on generative research synthesis, and our dataset task focuses on a complex research synthesis task derived from high-quality academic papers. Moreover, our automated evaluation approach is tailored to holistically assess the three key capabilities of generative research synthesis: retrieval, knowledge synthesis and verifiability.
AcademicEval (Zhang et al., 2024) evaluates long-context generation tasks using an ArXiv-derived dataset, similar to ours, however, its task differs substantially, focusing on summarization without retrieval, which is a key component of our task and of generative research synthesis systems more broadly. While the AcademicEval task provides a fixed set of references to a summarization system, our task requires a system to navigate the live web to collect information, and we evaluate this crucial capability of generative research synthesis.
Question Answering Benchmarks. Several recent works on generative research synthesis focus their evaluation on question answering (QA) benchmarks, which, unlike DeepScholar-bench, do not evaluate complex long-form research synthesis tasks and instead focus on short-form answers, whcih can be easily evaluated for correctness. These question answering benchmarks include SimpleQA (Wei et al., 2024), FRAMES (Krishna et al., 2025), GAIA (Mialon et al., 2023), as BrowserComp (Wei et al., 2025), WebWalkerQA (Wu et al., 2025) or others traditionally used to evaluate retrieval-augmented generation (Wadden et al., 2020; Jin et al., 2019; Yang et al., 2018; Joshi et al., 2017; Kwiatkowski et al., 2019; Ho et al., 2020; Trivedi et al., 2022; Lee et al., 2023). While these benchmarks involve a retrieval component, their synthesis task differs substantially from the generative research synthesis task we study. Specifically, each of these QA benchmarks focus on short-form questions with easily verifiable answers and straightforward notions correctness. In contrast, our benchmark provides a framework for studying complex, long-form research synthesis tasks, which lack an absolute notion of correctness and admit many possible reasonable answers.
7. Conclusion
In this work, we introduced DeepScholar-bench, a live dataset and holistic, automated evaluation framework designed to rigorously benchmark an emerging class of systems designed for generative research synthesis. By automatically sourcing queries from high-quality, recent ArXiv papers, our benchmark mitigates the risks of data staleness and training contamination, while offering a real research synthesis task. Moreover, DeepScholar-bench provides an automated evaluation to holistically measure three critical dimensions: retrieval quality, knowledge synthesis and verifiability. We further release DeepScholar-base, a reference pipeline, which we find provides a strong baseline for generative research synthesis. Overall our systematic evalaution of prior open-source systems, search AI’s, OpenAI’s DeepResearch and DeepScholar-base demonstrates significant opportunities for future work, with no system exceeding a score of across all metrics. These results demonstrate both the difficulty of DeepScholar-bench and the exciting opportunity for further advancement in this space. We hope that DeepScholar-bench and DeepScholar-base will support the development of more capable AI systems for generative research synthesis.
Acknowledgements.
This research was supported in part by affiliate members and other supporters of the Stanford DAWN project, including Meta, Google, and VMware, as well as Cisco, SAP, and a Sloan Fellowship. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.References
- (1)
- noa ([n.d.]a) [n.d.]a. arXiv.org e-Print archive. https://arxiv.org/
- noa ([n.d.]b) [n.d.]b. Claude takes research to new places. https://www.anthropic.com/news/research
- noa ([n.d.]c) [n.d.]c. Gemini Deep Research — your personal research assistant. https://gemini.google/overview/deep-research/
- noa ([n.d.]d) [n.d.]d. Gemini models | Gemini API. https://ai.google.dev/gemini-api/docs/models
- noa ([n.d.]e) [n.d.]e. Grok 3 Beta — The Age of Reasoning Agents | xAI. https://x.ai/news/grok-3
- noa ([n.d.]f) [n.d.]f. Introducing Claude 4. https://www.anthropic.com/news/claude-4
- noa ([n.d.]g) [n.d.]g. Introducing deep research | OpenAI. https://openai.com/index/introducing-deep-research/
- noa ([n.d.]h) [n.d.]h. Introducing Perplexity Deep Research. https://www.perplexity.ai/hub/blog/introducing-perplexity-deep-research
- noa ([n.d.]i) [n.d.]i. Model - OpenAI API. https://platform.openai.com
- noa ([n.d.]j) [n.d.]j. OpenAlex: The open catalog to the global research system | OpenAlex. https://openalex.org/
- noa ([n.d.]k) [n.d.]k. Su-Sea/ydc-deep-research-evals: you.com’s framework for evaluating deep research systems. https://github.com/Su-Sea/ydc-deep-research-evals
- noa ([n.d.]l) [n.d.]l. TREC CAR ← TREC CAR. https://trec-car.cs.unh.edu/
- noa (2025a) 2025a. Llama 4 - a meta-llama Collection. https://huggingface.co/collections/meta-llama/llama-4-67f0c30d9fe03840bc9d0164
- noa (2025b) 2025b. lotus-data/lotus. https://github.com/lotus-data/lotus original-date: 2024-07-16T16:39:06Z.
- noa (2025c) 2025c. meta-llama/Llama-4-Scout-17B-16E-Instruct · Hugging Face. https://huggingface.co/meta-llama/Llama-4-Scout-17B-16E-Instruct
- noa (2025d) 2025d. Wikipedia:Good article criteria. https://en.wikipedia.org/w/index.php?title=Wikipedia:Good_article_criteria&oldid=1303221295 Page Version ID: 1303221295.
- Alzubi et al. (2025) Salaheddin Alzubi, Creston Brooks, Purva Chiniya, Edoardo Contente, Chiara von Gerlach, Lucas Irwin, Yihan Jiang, Arda Kaz, Windsor Nguyen, Sewoong Oh, Himanshu Tyagi, and Pramod Viswanath. 2025. Open Deep Search: Democratizing Search with Open-source Reasoning Agents. https://arxiv.org/abs/2503.20201v1
- Arabzadeh and Clarke (2025) Negar Arabzadeh and Charles L. A. Clarke. 2025. Benchmarking LLM-based Relevance Judgment Methods. In Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (Padua, Italy) (SIGIR ’25). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3726302.3730305
- Asai et al. (2024) Akari Asai, Jacqueline He, Rulin Shao, Weijia Shi, Amanpreet Singh, Joseph Chee Chang, Kyle Lo, Luca Soldaini, Sergey Feldman, Mike D’arcy, David Wadden, Matt Latzke, Minyang Tian, Pan Ji, Shengyan Liu, Hao Tong, Bohao Wu, Yanyu Xiong, Luke Zettlemoyer, Graham Neubig, Dan Weld, Doug Downey, Wen-tau Yih, Pang Wei Koh, and Hannaneh Hajishirzi. 2024. OpenScholar: Synthesizing Scientific Literature with Retrieval-augmented LMs. https://doi.org/10.48550/arXiv.2411.14199 arXiv:2411.14199 [cs].
- Faggioli et al. (2023) Guglielmo Faggioli, Laura Dietz, Charles Clarke, Gianluca Demartini, Matthias Hagen, Claudia Hauff, Noriko Kando, Evangelos Kanoulas, Martin Potthast, Benno Stein, and Henning Wachsmuth. 2023. Perspectives on Large Language Models for Relevance Judgment. In Proceedings of the 2023 ACM SIGIR International Conference on Theory of Information Retrieval. 39–50. https://doi.org/10.1145/3578337.3605136 arXiv:2304.09161 [cs].
- Gao et al. (2023) Tianyu Gao, Howard Yen, Jiatong Yu, and Danqi Chen. 2023. Enabling Large Language Models to Generate Text with Citations. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 6465–6488. https://doi.org/10.18653/v1/2023.emnlp-main.398
- Ho et al. (2020) Xanh Ho, Anh-Khoa Duong Nguyen, Saku Sugawara, and Akiko Aizawa. 2020. Constructing A Multi-hop QA Dataset for Comprehensive Evaluation of Reasoning Steps. In Proceedings of the 28th International Conference on Computational Linguistics, Donia Scott, Nuria Bel, and Chengqing Zong (Eds.). International Committee on Computational Linguistics, Barcelona, Spain (Online), 6609–6625. https://doi.org/10.18653/v1/2020.coling-main.580
- Jin et al. (2019) Qiao Jin, Bhuwan Dhingra, Zhengping Liu, William Cohen, and Xinghua Lu. 2019. PubMedQA: A Dataset for Biomedical Research Question Answering. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, Hong Kong, China, 2567–2577. https://doi.org/10.18653/v1/D19-1259
- Joshi et al. (2017) Mandar Joshi, Eunsol Choi, Daniel Weld, and Luke Zettlemoyer. 2017. TriviaQA: A Large Scale Distantly Supervised Challenge Dataset for Reading Comprehension. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Regina Barzilay and Min-Yen Kan (Eds.). Association for Computational Linguistics, Vancouver, Canada, 1601–1611. https://doi.org/10.18653/v1/P17-1147
- Krishna et al. (2025) Satyapriya Krishna, Kalpesh Krishna, Anhad Mohananey, Steven Schwarcz, Adam Stambler, Shyam Upadhyay, and Manaal Faruqui. 2025. Fact, Fetch, and Reason: A Unified Evaluation of Retrieval-Augmented Generation. https://doi.org/10.48550/arXiv.2409.12941 arXiv:2409.12941 [cs].
- Kwiatkowski et al. (2019) Tom Kwiatkowski, Jennimaria Palomaki, Olivia Redfield, Michael Collins, Ankur Parikh, Chris Alberti, Danielle Epstein, Illia Polosukhin, Jacob Devlin, Kenton Lee, Kristina Toutanova, Llion Jones, Matthew Kelcey, Ming-Wei Chang, Andrew M. Dai, Jakob Uszkoreit, Quoc Le, and Slav Petrov. 2019. Natural Questions: A Benchmark for Question Answering Research. Transactions of the Association for Computational Linguistics 7 (2019), 452–466. https://doi.org/10.1162/tacl_a_00276 Place: Cambridge, MA Publisher: MIT Press.
- Kwon et al. (2023) Woosuk Kwon, Zhuohan Li, Siyuan Zhuang, Ying Sheng, Lianmin Zheng, Cody Hao Yu, Joseph E. Gonzalez, Hao Zhang, and Ion Stoica. 2023. Efficient Memory Management for Large Language Model Serving with PagedAttention. https://doi.org/10.48550/arXiv.2309.06180 arXiv:2309.06180 [cs].
- Lee et al. (2023) Yoonjoo Lee, Kyungjae Lee, Sunghyun Park, Dasol Hwang, Jaehyeon Kim, Hong-In Lee, and Moontae Lee. 2023. QASA: Advanced Question Answering on Scientific Articles. In Proceedings of the 40th International Conference on Machine Learning. PMLR, 19036–19052. https://proceedings.mlr.press/v202/lee23n.html
- Li et al. (2025) Dawei Li, Bohan Jiang, Liangjie Huang, Alimohammad Beigi, Chengshuai Zhao, Zhen Tan, Amrita Bhattacharjee, Yuxuan Jiang, Canyu Chen, Tianhao Wu, Kai Shu, Lu Cheng, and Huan Liu. 2025. From Generation to Judgment: Opportunities and Challenges of LLM-as-a-judge. https://doi.org/10.48550/arXiv.2411.16594 arXiv:2411.16594 [cs].
- Li et al. (2024) Ruosen Li, Teerth Patel, and Xinya Du. 2024. PRD: Peer Rank and Discussion Improve Large Language Model based Evaluations. https://doi.org/10.48550/arXiv.2307.02762 arXiv:2307.02762 [cs].
- Liu et al. (2023) Nelson Liu, Tianyi Zhang, and Percy Liang. 2023. Evaluating Verifiability in Generative Search Engines. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 7001–7025. https://doi.org/10.18653/v1/2023.findings-emnlp.467
- Mialon et al. (2023) Grégoire Mialon, Clémentine Fourrier, Craig Swift, Thomas Wolf, Yann LeCun, and Thomas Scialom. 2023. GAIA: a benchmark for General AI Assistants. https://doi.org/10.48550/arXiv.2311.12983 arXiv:2311.12983 [cs].
- Nanni et al. (2017) Federico Nanni, Bhaskar Mitra, Matt Magnusson, and Laura Dietz. 2017. Benchmark for Complex Answer Retrieval. https://doi.org/10.48550/arXiv.1705.04803 arXiv:1705.04803 [cs].
- Patel et al. (2025) Liana Patel, Siddharth Jha, Melissa Pan, Harshit Gupta, Parth Asawa, Carlos Guestrin, and Matei Zaharia. 2025. Semantic Operators: A Declarative Model for Rich, AI-based Data Processing. https://doi.org/10.48550/arXiv.2407.11418 arXiv:2407.11418 [cs].
- Pradeep et al. (2025) Ronak Pradeep, Nandan Thakur, Shivani Upadhyay, Daniel Campos, Nick Craswell, and Jimmy Lin. 2025. The Great Nugget Recall: Automating Fact Extraction and RAG Evaluation with Large Language Models. https://doi.org/10.48550/arXiv.2504.15068 arXiv:2504.15068 [cs].
- Qwen et al. (2025) Qwen, An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, Huan Lin, Jian Yang, Jianhong Tu, Jianwei Zhang, Jianxin Yang, Jiaxi Yang, Jingren Zhou, Junyang Lin, Kai Dang, Keming Lu, Keqin Bao, Kexin Yang, Le Yu, Mei Li, Mingfeng Xue, Pei Zhang, Qin Zhu, Rui Men, Runji Lin, Tianhao Li, Tianyi Tang, Tingyu Xia, Xingzhang Ren, Xuancheng Ren, Yang Fan, Yang Su, Yichang Zhang, Yu Wan, Yuqiong Liu, Zeyu Cui, Zhenru Zhang, and Zihan Qiu. 2025. Qwen2.5 Technical Report. https://doi.org/10.48550/arXiv.2412.15115 arXiv:2412.15115 [cs].
- Rahmani et al. (2024a) Hossein A. Rahmani, Clemencia Siro, Mohammad Aliannejadi, Nick Craswell, Charles L. A. Clarke, Guglielmo Faggioli, Bhaskar Mitra, Paul Thomas, and Emine Yilmaz. 2024a. LLM4Eval: Large Language Model for Evaluation in IR. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval. ACM, Washington DC USA, 3040–3043. https://doi.org/10.1145/3626772.3657992
- Rahmani et al. (2024b) Hossein A. Rahmani, Emine Yilmaz, Nick Craswell, Bhaskar Mitra, Paul Thomas, Charles L. A. Clarke, Mohammad Aliannejadi, Clemencia Siro, and Guglielmo Faggioli. 2024b. LLMJudge: LLMs for Relevance Judgments. https://doi.org/10.48550/arXiv.2408.08896 arXiv:2408.08896 [cs].
- Shao et al. (2024) Yijia Shao, Yucheng Jiang, Theodore A. Kanell, Peter Xu, Omar Khattab, and Monica S. Lam. 2024. Assisting in Writing Wikipedia-like Articles From Scratch with Large Language Models. https://doi.org/10.48550/arXiv.2402.14207 arXiv:2402.14207 [cs].
- Soldaini et al. (2024) Luca Soldaini, Rodney Kinney, Akshita Bhagia, Dustin Schwenk, David Atkinson, Russell Authur, Ben Bogin, Khyathi Chandu, Jennifer Dumas, Yanai Elazar, Valentin Hofmann, Ananya Jha, Sachin Kumar, Li Lucy, Xinxi Lyu, Nathan Lambert, Ian Magnusson, Jacob Morrison, Niklas Muennighoff, Aakanksha Naik, Crystal Nam, Matthew Peters, Abhilasha Ravichander, Kyle Richardson, Zejiang Shen, Emma Strubell, Nishant Subramani, Oyvind Tafjord, Evan Walsh, Luke Zettlemoyer, Noah Smith, Hannaneh Hajishirzi, Iz Beltagy, Dirk Groeneveld, Jesse Dodge, and Kyle Lo. 2024. Dolma: an Open Corpus of Three Trillion Tokens for Language Model Pretraining Research. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 15725–15788. https://doi.org/10.18653/v1/2024.acl-long.840
- Thakur et al. (2021) Nandan Thakur, Nils Reimers, Andreas Rücklé, Abhishek Srivastava, and Iryna Gurevych. 2021. BEIR: A Heterogenous Benchmark for Zero-shot Evaluation of Information Retrieval Models. https://doi.org/10.48550/arXiv.2104.08663 arXiv:2104.08663 [cs].
- Thomas et al. (2024) Paul Thomas, Seth Spielman, Nick Craswell, and Bhaskar Mitra. 2024. Large language models can accurately predict searcher preferences. https://doi.org/10.48550/arXiv.2309.10621 arXiv:2309.10621 [cs].
- Trivedi et al. (2022) Harsh Trivedi, Niranjan Balasubramanian, Tushar Khot, and Ashish Sabharwal. 2022. MuSiQue: Multihop Questions via Single-hop Question Composition. https://doi.org/10.48550/arXiv.2108.00573 arXiv:2108.00573 [cs].
- Upadhyay et al. (2024a) Shivani Upadhyay, Ronak Pradeep, Nandan Thakur, Daniel Campos, Nick Craswell, Ian Soboroff, Hoa Trang Dang, and Jimmy Lin. 2024a. A Large-Scale Study of Relevance Assessments with Large Language Models: An Initial Look. https://doi.org/10.48550/arXiv.2411.08275 arXiv:2411.08275 [cs].
- Upadhyay et al. (2024b) Shivani Upadhyay, Ronak Pradeep, Nandan Thakur, Nick Craswell, and Jimmy Lin. 2024b. UMBRELA: UMbrela is the (Open-Source Reproduction of the) Bing RELevance Assessor. https://doi.org/10.48550/arXiv.2406.06519 arXiv:2406.06519 [cs].
- Voorhees (2009) Ellen M. Voorhees. 2009. I Come Not To Bury Cranfield, but to Praise It. NIST (Oct. 2009). https://www.nist.gov/publications/i-come-not-bury-cranfield-praise-it Last Modified: 2017-02-19T20:02-05:00 Publisher: Ellen M. Voorhees.
- Wadden et al. (2020) David Wadden, Shanchuan Lin, Kyle Lo, Lucy Lu Wang, Madeleine van Zuylen, Arman Cohan, and Hannaneh Hajishirzi. 2020. Fact or Fiction: Verifying Scientific Claims. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 7534–7550. https://doi.org/10.18653/v1/2020.emnlp-main.609
- Wei et al. (2024) Jason Wei, Nguyen Karina, Hyung Won Chung, Yunxin Joy Jiao, Spencer Papay, Amelia Glaese, John Schulman, and William Fedus. 2024. Measuring short-form factuality in large language models. https://doi.org/10.48550/arXiv.2411.04368 arXiv:2411.04368 [cs].
- Wei et al. (2025) Jason Wei, Zhiqing Sun, Spencer Papay, Scott McKinney, Jeffrey Han, Isa Fulford, Hyung Won Chung, Alex Tachard Passos, William Fedus, and Amelia Glaese. 2025. BrowseComp: A Simple Yet Challenging Benchmark for Browsing Agents. https://doi.org/10.48550/arXiv.2504.12516 arXiv:2504.12516 [cs].
- Worledge et al. (2024) Theodora Worledge, Tatsunori Hashimoto, and Carlos Guestrin. 2024. The Extractive-Abstractive Spectrum: Uncovering Verifiability Trade-offs in LLM Generations. https://doi.org/10.48550/arXiv.2411.17375 arXiv:2411.17375 [cs].
- Wu et al. (2025) Jialong Wu, Wenbiao Yin, Yong Jiang, Zhenglin Wang, Zekun Xi, Runnan Fang, Linhai Zhang, Yulan He, Deyu Zhou, Pengjun Xie, and Fei Huang. 2025. WebWalker: Benchmarking LLMs in Web Traversal. https://doi.org/10.48550/arXiv.2501.07572 arXiv:2501.07572 [cs].
- Yang et al. (2018) Zhilin Yang, Peng Qi, Saizheng Zhang, Yoshua Bengio, William W. Cohen, Ruslan Salakhutdinov, and Christopher D. Manning. 2018. HotpotQA: A Dataset for Diverse, Explainable Multi-hop Question Answering. https://doi.org/10.48550/arXiv.1809.09600 arXiv:1809.09600 [cs].
- Zhang et al. (2024) Haozhen Zhang, Tao Feng, Pengrui Han, and Jiaxuan You. 2024. AcademicEval: Live Long-Context LLM Benchmark. (Oct. 2024). https://openreview.net/forum?id=iRYExPKnxm
- Zheng et al. (2025) Yuxiang Zheng, Dayuan Fu, Xiangkun Hu, Xiaojie Cai, Lyumanshan Ye, Pengrui Lu, and Pengfei Liu. 2025. DeepResearcher: Scaling Deep Research via Reinforcement Learning in Real-world Environments. https://arxiv.org/abs/2504.03160v4
- Zheng et al. (2024) Yuxiang Zheng, Shichao Sun, Lin Qiu, Dongyu Ru, Cheng Jiayang, Xuefeng Li, Jifan Lin, Binjie Wang, Yun Luo, Renjie Pan, Yang Xu, Qingkai Min, Zizhao Zhang, Yiwen Wang, Wenjie Li, and Pengfei Liu. 2024. OpenResearcher: Unleashing AI for Accelerated Scientific Research. In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, Delia Irazu Hernandez Farias, Tom Hope, and Manling Li (Eds.). Association for Computational Linguistics, Miami, Florida, USA, 209–218. https://doi.org/10.18653/v1/2024.emnlp-demo.22
8. Appendix
8.1. DeepScholar-bench Dataset
We provide a detailed overview and schema of the DeepScholar-bench dataset in Figure 4.
8.2. Baselines

### v2

@@ A.3.5 Manual Validation of LLM-based Evaluation
We study the alignment between our LLM-based evaluation and human judgments to assess the effectiveness of our automated metrics. Overall, we find that each of the metrics we introduce for the DeepScholar-bench task exhibit high agreement between LLM-based judgments and human annotations. We collect over 400 human annotations, and Table 10 shows the agreement score between human and LLMs for each LLM classification task associated with each automated metrics. The results demonstrate above agreement scores across each. We additionally compute the nugget precision and grounded-ness scores, following prior work [46], observing scores of and respectively, demonstrating that the generated nuggets are accurate and do not contain hallucinations.
A.3.6 Agreement Between Different LLM Judges
To test the robustness of our evaluation to the choice of LLM judge, we repeat all experiments with three different judges: GPT-4o, Llama-4-17b-16e-Instruct, and DeepSeek-R1-Distill-Qwen-32B. For each pair of judges, we compute the Pearson correlation between system-level scores across five baselines (DeepScholar-ref (GPT-4.1, Claude), DeepScholar-ref (Llama-4), Search Agent (Claude), Search Agent (Llama-4), and OpenAI DeepResearch for all metrics except document importance, which is not LLM-based. We then run perturbation-based permutation tests to assess whether these correlations are significantly greater than zero. The resulting correlations and significance markers are shown in Table 11.
Correlations are generally very high (often above 0.9). Notably GPT-4o vs. DeepSeek shows statistically significant correlations across all the metrics at the level, indicating that these two judges high agreement. Most metrics also remain significantly correlated, but we observe the main sensitivity to the choice of judge on nugget coverage (non-significant for both GPT-4o vs. Llama-4-17b-16e-Instruct and DeepSeek-R1-Distill-Qwen-32B vs. Llama-4-17b-16e-Instruct). In addition, relevance rate shows non-statistically significant correlation for GPT-4o vs. Llama-4-17b-16e-Instruct and similarly Citation Precision for DeepSeek-R1-Distill-Qwen-32B vs. Llama-4-17b-16e-Instruct. It is worth mentioning that even in these cases, the correlations remain large (all above 0.41), suggesting that discrepancies are not severe.
| Judge pair | Org. | Nug. | Cov. Rel. Rate | Ref. Cov. | Cite-P | Claim Cov () |
| GPT-4o vs. Llama-4 | 0.985∗ | 0.413 | 0.941 | 0.975∗ | 0.817∗ | 0.958∗ |
| GPT-4o vs. DeepSeek | 0.966∗ | 0.920∗ | 0.980∗ | 0.990∗ | 0.843∗ | 0.996∗ |
| DeepSeek vs. Llama-4 | 0.991∗ | 0.668 | 0.942∗ | 0.994∗ | 0.986 | 0.963∗ |
A.4 Evaluation Details
Here we provide the detailed implementation and prompts for the evaluation metrics used in DeepScholar-Bench. In the following, we also run some ablation studies on different metrics.
Knowledge Synthesis – Organization: Box A.4.3 shows the prompt used to compare system-generated reports with the human-written exemplar, yielding the win-rate of each system based on organization and coherence.
Knowledge Synthesis – Nugget Coverage: For this metric we follow the nugget-based evaluation prompts from prior work [39], which extract essential facts from the exemplar and check their presence in the generated report.
Retrieval Quality – Document Importance: We select important references using the LOTUS program shown in Figure 5.
Retrieval Quality – Relevance Rate: Following the Cranfield model of relevance, in Box A.4.3 we prompt the model for assigning graded relevance scores (0–2) to each retrieved source.
We note that Reference Coverage and Document Importance are computed deterministically as explained in Section 3.2: the former by comparing the retrieved set against the important references in the human-written exemplar, and the latter by normalizing the number of citations of the reference papers.
A.4.1 Reference Coverage
A.4.2 Ablation Study on Verifiability
A.4.3 Document Importance Across Human Exemplars
In this section, we illustrate the distribution of document importance, measured by the number of citations of references in the human-written exemplars in DeepScholar-Bench. Figure 7 reports two histograms: (a) the distribution of citation counts across all references, and (b) the distribution restricted to references that appear on ArXiv. We plot the logarithm of citation counts, with values obtained from the OpenAlex API [35], an open and widely used scholarly database that provides citation-level metadata. While citation counts in OpenAlex may not exactly match those from other sources such as Google Scholar, the relative counts are consistent, making it a reliable open-source alternative.
A.5 Extended Description of Baselines